OpenStudy (anonymous):

Find the points of intersection of the graphs of the equations.(2 separate questions) 1.(x^2)+(y^2)=25 2x+y=10 2.(x^2)+(y^2)=5 x-y=1

OpenStudy (tkhunny):

Have you considered substitution? 1) \(x^{2} + y^{2} = 25\) 2) \(2x + y = 10 \implies y = 10-2x\;and\;y^{2} = (10 - 2x)^{2}\) Substituting 2) into 1) \(x^{2} + (10 - 2x)^{2} = 25\) Expand, simplify where possible, and solve.